# The Ranking-Selberg integral on ${\bf GSp}(2)$ for square free levels

### v1

@@ 6.5. Computation for newforms of type ${\rm III}a$ (case5)
We now consider the values of $B_{1}$ and $B_{2}$ at the identity element. We remark that [32, Lemma 9.2] ensures its non-vanishing and
$\Lambda(\varpi)=\alpha\gamma^{2}$
by the condition of central character of $\pi$.
@@
Let $S\in{\rm{Sym^{2}}}(F)$ such that $L/F$ is an unramified field extension. Suppose that $\mu:F^{\times}\to{\mathbb{C}}^{1}$ and $\Lambda:L^{\times}\to{\mathbb{C}}^{1}$ are unramified characters. Then we have
@@
For $B\in{\mathscr{B}}(S,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}$ and any non-negative integer $l$, by Lemmas 5.1 and 5.3 in [32], we get the following relations;
@@
From (6.41), (6.42), (6.43), and (6.44), we obtain $B_{2}(1_{4})=\alpha^{-1}B_{1}(1_{4})$. The non-vanishing of $B_{1}$ at $1_{4}$ is a consequence of [32, Theorem 9.3]. ∎
@@
Suppose that $\pi$ is of type IIIa. Let $S\in{\rm{Sym^{2}}}(F)$ be such that $L/F$ is an unramified field extension. Suppose that $\mu:F^{\times}\to{\mathbb{C}}^{1}$ and $\Lambda:L^{\times}\to{\mathbb{C}}^{1}$ are unramified characters. When $\phi=1_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}$ and ${\operatorname{Re}}(s)>1$, we have
@@
By substituting $B=B_{i}$ in (6.34) for $i=1,2$, the statement is immediate for $B=B_{i}$ by (6.39), the first relation in (6.40), and Lemma 6.9. Since $B_{1}$ and $B_{2}$ span ${\mathscr{B}}(S,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}$, we complete the proof. ∎
@@
Let $\pi$, $S$, $\mu$, and $\Lambda$ be as in Proposition 6.10, then
@@
6.6. Computation for newforms of type ${\rm VI}b$ (case6)
Suppose that $\pi$ is of type VIb. By [32, Table 3], any element $B\in{\mathscr{B}}(S,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}$ satisfies
@@ 6.6. Computation for newforms of type ${\rm VI}b$ (case6)
By [32, Theorem 9.3], the value $B(1_{4})\not=0$ if and only if $B=0$.
@@
Suppose that $\pi$ is of type VIb. Let $S\in{\rm{Sym^{2}}}(F)$ be such that $L/F$ is an unramified field extension. Suppose that $\mu:F^{\times}\to{\mathbb{C}}^{1}$ and $\Lambda:L^{\times}\to{\mathbb{C}}^{1}$ are unramified characters. When $\phi=1_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}$ and ${\operatorname{Re}}(s)>1$, we have
@@
In our case, the relations (6.40) are also true. The proof is similar to that of Proposition 6.10. ∎
@@
Let $\pi$, $S$, $\mu$, and $\Lambda$ be as in Proposition 6.12, then
@@ 6.7. The values of Bessel models at $1_{4}$ for old forms
By [33, Table 3], the space $V^{I}$ is 8-dimensional space spanned by
@@
The representation matrices of operators $T_{1,0}$ and $\eta$ on $V_{\pi}^{K_{0}({\mathfrak{p}})}$ are given in [33, Table 3] as follows;
(6.49) $\begin{array}[]{|c|c|c|}\hline\cr{}\hfil&{{\rm type\ I}}&{{\rm type\ IIb}}\\ \hline\cr{T_{1,0}}&{\left[\begin{smallmatrix}\alpha\beta\gamma q^{3/2}&0&0&0\\ \alpha\beta\gamma(q-1)q^{1/2}&\beta\gamma q^{3/2}&0&0\\ \alpha\beta\gamma(q-1)q^{1/2}&\beta\gamma(q-1)q^{1/2}&\alpha\gamma q^{3/2}&0\\ \alpha\beta\gamma(q-1)q^{1/2}&\beta\gamma(q-1)q^{1/2}&\alpha\gamma(q-1)q^{1/2}&\gamma q^{3/2}\end{smallmatrix}\right]}&{\left[\begin{smallmatrix}\alpha^{2}\gamma q^{3/2}&0&0\\ \alpha^{2}\gamma(q-1)q^{1/2}&\alpha\gamma q^{2}&0\\ \alpha^{2}\gamma(q-1)q^{1/2}&\alpha\gamma(q^{2}-1)&\gamma q^{3/2}\end{smallmatrix}\right]}\\ \hline\cr{\eta}&{\left[\begin{smallmatrix}0&0&0&\gamma q^{3/2}\\ 0&0&\alpha\gamma q^{1/2}&0\\ 0&\beta\gamma q^{-1/2}&0&0\\ \alpha\beta\gamma q^{-3/2}&0&0&0\end{smallmatrix}\right]}&{\left[\begin{smallmatrix}0&0&\gamma q^{3/2}\\ 0&\alpha\gamma&0\\ \alpha^{2}\gamma q^{-3/2}&0&0\end{smallmatrix}\right]}\\ \hline\cr\end{array}$
Now we assume that $\pi$ has the trivial character. Let $(\ell_{\pi}^{0},f_{K})$ be the Bessel data for $\pi$ defined above. For $\bullet\in\{{\rm I},{\rm IIb}\}$, we set a basis $\{B_{i}^{\bullet}\}_{1{\leqslant}i{\leqslant}\dim V_{\pi}^{K_{0}({\mathfrak{p}})}}$ of $V_{\pi}^{K_{0}({\mathfrak{p}})}$ by
@@
Then the following result holds.
@@
Set $d:=\dim V_{\pi}^{K_{0}({\mathfrak{p}})}$. By the construction of the elements $\{f_{i}^{\bullet}\}_{1{\leqslant}i{\leqslant}d}$, we have $B_{\pi}^{0}=\sum_{i=1}^{d}B_{i}^{\bullet}$. For any $\ell\in{\mathbb{Z}}_{{\geqslant}0}$, we act by $T_{1,0}^{l}$ on the both sides and evaluate for $g=1_{4}$, then the first relation in (6.40) implies that
$B_{\pi}^{0}(h(\ell,0))=\sum_{i=1}^{d}(T_{1,0}^{\ell}B_{i}^{\bullet})(1_{4}).$
The values of the left hand side is given in [6, Corollary 1.8]. Hence by using the table (6.49) and a direct calculation, we can solve the values of $B_{i}^{\bullet}(1_{4})$. ∎
@@ 6.8. Local periods
$\mathbb{I}_{\pi,K_{0}({\mathfrak{p}})}^{(s)}(\xi^{\pi},\phi,\Lambda,\mu;g)=\sum_{B\in\mathcal{B}((V_{\pi}^{*})^{S,\Lambda},K_{0}({\mathfrak{p}}))}Z^{*}(\phi,B,s,\mu;g)\overline{B(1_{4})}.$
@@
is the normalized zeta integral and $\mathcal{B}((V_{\pi}^{*})^{S,\Lambda},K_{0}({\mathfrak{p}}))$ is an orthonormal basis of ${\mathscr{B}}(S,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}$ with respect to the inner product on ${\mathscr{B}}(S,\Lambda)[\pi]$ via the $\mathsf{G}(F)$-isomorphism
(6.50) $V_{\pi}\owns\xi\mapsto(g\mapsto\ell_{\pi}(\pi(g)\xi))\in{\mathscr{B}}(S,\Lambda)[\pi].$
@@ 6.8.1. Local periods for type I and IIb
For $\bullet\in\{{\rm I},{\rm IIb}\}$, the set $\{\langle B_{i}^{\bullet}|B_{i}^{\bullet}\rangle^{-\frac{1}{2}}B_{i}^{\bullet}\}_{1{\leqslant}i{\leqslant}\dim{V_{\pi}^{K_{0}({\mathfrak{p}})}}}$ forms an orthonormal basis of ${\mathscr{B}}(S,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}$. By [37], the values $\langle B_{i}^{\bullet}|B_{i}^{\bullet}\rangle$ are given as follows:
@@
By Propositions 6.8 and 6.14, and the matrix representation of $T_{1,0}$ and $\eta$ given in (6.49), we can calculate the local period by a direct calculation.  [^2] We have used Mathematica 13.
@@ 6.8.2. Local period for type IIIa
Suppose that $\pi$ is of type IIIa. Then, as remarked before, $|\alpha|=|\gamma|=1$ and $\alpha\not=1$. By Lemma 6.9, there exists a basis $\{B_{i}\}_{i=1,2}$ of ${\mathscr{B}}(S,\Lambda)[\pi]^{K_{(}{\mathfrak{p}})}$, unique up to constant, satisfying the relations (6.39) and $B_{2}(1_{4})=\alpha^{-1}B_{1}(1_{4})\not=0$. Thus, we can uniquely fix $\{B_{i}^{{\rm IIIa}}\}_{i=1,2}$ by requiring $B_{1}^{{\rm IIIa}}(1_{4})=1$ and the datum $(\ell_{\pi},\xi^{\pi})$ by $\ell_{\pi}(\xi^{\pi})=B^{{\rm IIIa}}_{1}$ to form (6.50) and induce the $\mathsf{G}(F)$-invariant inner product on ${\mathscr{B}}(S,\Lambda)[\pi]$ such that $\langle B_{1}^{{\rm IIIa}}|B_{1}^{{\rm IIIa}}\rangle=1$. Since $\langle B_{1}|B_{2}\rangle=0$ by [7, Lemma 2.11], we then have $\langle B_{-}^{{\rm IIIa}}|B_{-}^{{\rm IIIa}}\rangle=1$ and $\langle B_{+}^{{\rm IIIa}}|B_{-}^{{\rm IIIa}}\rangle=0$ due to $|\alpha|=1$.
@@
Let $\pi$ be a smooth admissible irreducible unitary representation of type IIIa. Suppose that $:F^{\times}\to{\mathbb{C}}^{1}$ is an unramified character. When $\phi=\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}$, we have
$\mathbb{I}_{\pi,K_{0}({\mathfrak{p}})}^{(s)}(\xi^{\pi},\phi,\Lambda,\mu;\eta)=\frac{2\Lambda(\varpi)^{-1}\mu(\varpi)^{-1}q^{s+1}}{q^{2}+1}$

### v2

@@ 6.6. Computation for newforms of type ${\rm III}a$ (case 7)
We now consider the values of $B_{1}$ and $B_{2}$ at the identity element. We remark that [33, Lemma 9.2] ensures its non-vanishing and $\Lambda(\varpi)=\alpha\gamma^{2}$ by the condition of central character of $\pi$.
@@
Let $T\in{\rm{Sym^{2}}}(F)$ such that $L/F$ is an unramified field extension. Suppose that $\mu:F^{\times}\to{\mathbb{C}}^{1}$ and $\Lambda:L^{\times}\to{\mathbb{C}}^{1}$ are unramified characters. Then we have
@@
For $B\in{\mathscr{B}}(T,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}$ and any non-negative integer $l$, by Lemmas 5.1 and 5.3 in [33], we get the following relations;
@@
From (6.35), (6.36), (6.37), and (6.38), we obtain $B_{2}(1_{4})=\alpha^{-1}B_{1}(1_{4})$. The non-vanishing of $B_{1}$ at $1_{4}$ is a consequence of [33, Theorem 9.3]. ∎
@@
Suppose that $\pi$ is of type IIIa. Let $T\in{\rm{Sym^{2}}}(F)$ be such that $L/F$ is an unramified field extension. Suppose that $\mu:F^{\times}\to{\mathbb{C}}^{1}$ and $\Lambda:L^{\times}\to{\mathbb{C}}^{1}$ are unramified characters. When $\phi=1_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}$ and ${\operatorname{Re}}(s)>1$, we have
@@
By substituting $B=B_{i}$ in (6.28) for $i=1,2$, the statement is immediate for $B=B_{i}$ by (6.33), the first relation in (6.34), and Lemma 6.10. Since $B_{1}$ and $B_{2}$ span ${\mathscr{B}}(T,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}$, we complete the proof. ∎
@@
Let $\pi$, $T$, $\mu$, and $\Lambda$ be as in Proposition 6.11, then
@@
6.7. Computation for newforms of type ${\rm VI}b$ (case 8)
Suppose that $\pi$ is of type VIb. By [33, Table 3], any element $B\in{\mathscr{B}}(T,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}$ satisfies
@@ 6.7. Computation for newforms of type ${\rm VI}b$ (case 8)
By [33, Theorem 9.3], the value $B(1_{4})\not=0$ if and only if $B=0$.
@@
Suppose that $\pi$ is of type VIb. Let $T\in{\rm{Sym^{2}}}(F)$ be such that $L/F$ is an unramified field extension. Suppose that $\mu:F^{\times}\to{\mathbb{C}}^{1}$ and $\Lambda:L^{\times}\to{\mathbb{C}}^{1}$ are unramified characters. When $\phi=1_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}$ and ${\operatorname{Re}}(s)>1$, we have
$Z(\phi,B,s,\mu;\eta)=\frac{\Lambda(\varpi)^{-1}\mu(\varpi)^{-1}q^{s+1}}{q^{2}+1}L(s+1/2,\pi,\mu)\cdot B(1_{4}),\quad B\in{\mathscr{B}}(T,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}.$
@@
The proof is similar to that of Proposition 6.11 by using relations in (6.34). ∎
@@
Let $\pi$, $T$, $\mu$, and $\Lambda$ be as in Proposition 6.13, then
@@ 6.8. The values of Bessel models at $1_{4}$ for old forms
By [34, Table 3], the space $V^{I}$ is 8-dimensional space spanned by
@@
The representation matrices of operators $T_{1,0}$ and $\eta$ on $V_{\pi}^{K_{0}({\mathfrak{p}})}$ are given in [34, Table 3]. Now we assume that $\pi$ has the trivial character. Let $(\ell_{\pi}^{0},f_{K})$ be the Bessel data for $\pi$ defined above. For $\bullet\in\{{\rm I},{\rm IIb}\}$, define a basis $\{B_{i}^{\bullet}\}_{1{\leqslant}i{\leqslant}\dim V_{\pi}^{K_{0}({\mathfrak{p}})}}$ of $V_{\pi}^{K_{0}({\mathfrak{p}})}$ by
@@
We need the values of these functions at $g=1_{4}$, which are given as follows.
@@
Set $d:=\dim V_{\pi}^{K_{0}({\mathfrak{p}})}$. By the construction of the elements $\{f_{i}^{\bullet}\}_{1{\leqslant}i{\leqslant}d}$, we have $B_{\pi}^{0}=\sum_{i=1}^{d}B_{i}^{\bullet}$. For any $\ell\in{\mathbb{Z}}_{{\geqslant}0}$, we let $T_{1,0}^{l}$ act on the both sides of this equation and then evaluate at $g=1_{4}$ using the first relation in (6.34); thus,
$B_{\pi}^{0}(h(\ell,0))=q^{-3\ell}\sum_{i=1}^{d}(T_{1,0}^{\ell}B_{i}^{\bullet})(1_{4}).$
The values of the left hand side is given in [6, Corollary 1.8]; by [34, Table 3], we have a system of linear equations among $B_{i}^{\bullet}(1_{4})$, which is solved easily. ∎
@@ 6.9. Local periods
$\mathbb{I}_{\pi,K_{0}({\mathfrak{p}})}^{(s)}(\xi^{\pi},\phi,\Lambda,\mu;g)=\sum_{B\in\mathcal{B}((V_{\pi}^{*})^{T,\Lambda},K_{0}({\mathfrak{p}}))}Z^{*}(\phi,B,s,\mu;g)\overline{B(1_{4})},$
@@
is the normalized zeta integral and $\mathcal{B}((V_{\pi}^{*})^{T,\Lambda},K_{0}({\mathfrak{p}}))$ is an orthonormal basis of ${\mathscr{B}}(T,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}$ with respect to the inner product on ${\mathscr{B}}(T,\Lambda)[\pi]$ via the $\mathsf{G}(F)$-isomorphism
(6.40) $V_{\pi}\owns\xi\mapsto(g\mapsto\ell_{\pi}(\pi(g)\xi))\in{\mathscr{B}}(T,\Lambda)[\pi].$
@@ 6.9.1. Local periods for type I and IIb
For $\bullet\in\{{\rm I},{\rm IIb}\}$, the set $\{\langle B_{i}^{\bullet}|B_{i}^{\bullet}\rangle^{-\frac{1}{2}}B_{i}^{\bullet}\}_{1{\leqslant}i{\leqslant}\dim{V_{\pi}^{K_{0}({\mathfrak{p}})}}}$ forms an orthonormal basis of ${\mathscr{B}}(T,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}$. By [38], the values $\langle B_{i}^{\bullet}|B_{i}^{\bullet}\rangle$ are given as follows:
@@
By Propositions 6.9 and 6.15, and the matrix representation of $T_{1,0}$ and $\eta$ given in [34, Table 3], we can calculate the local period by a direct calculation.  [^2] We have used Mathematica 13.
@@ 6.9.2. Local period for type IIIa
Suppose that $\pi$ is of type IIIa. Then, as remarked before, $|\alpha|=|\gamma|=1$ and $\alpha\not=1$. By Lemma 6.10, there exists a basis $\{B_{i}\}_{i=1,2}$ of ${\mathscr{B}}(T,\Lambda)[\pi]^{K_{(}{\mathfrak{p}})}$, unique up to constant, satisfying the relations (6.33) and $B_{2}(1_{4})=\alpha^{-1}B_{1}(1_{4})\not=0$. Thus, we can uniquely fix $\{B_{i}^{{\rm IIIa}}\}_{i=1,2}$ by requiring $B_{1}^{{\rm IIIa}}(1_{4})=1$ and the datum $(\ell_{\pi},\xi^{\pi})$ by $\ell_{\pi}(\xi^{\pi})=B^{{\rm IIIa}}_{1}$ to form (6.40) and induce the $\mathsf{G}(F)$-invariant inner product on ${\mathscr{B}}(T,\Lambda)[\pi]$ such that $\langle B_{1}^{{\rm IIIa}}|B_{1}^{{\rm IIIa}}\rangle=1$. Since $\langle B_{1}|B_{2}\rangle=0$ by [7, Lemma 2.11], we then have $\langle B_{2}^{{\rm IIIa}}|B_{2}^{{\rm IIIa}}\rangle=1$ and $\langle B_{1}^{{\rm IIIa}}|B_{2}^{{\rm IIIa}}\rangle=0$ due to $|\alpha|=1$. The following proposition is immediate from Proposition 6.11.
@@
Let $\pi$ be a smooth admissible irreducible unitary representation of type IIIa. Suppose that $:F^{\times}\to{\mathbb{C}}^{1}$ is an unramified character. When $\phi=\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}$,
$\mathbb{I}_{\pi,K_{0}({\mathfrak{p}})}^{(s)}(\xi^{\pi},\phi,\Lambda,\mu;\eta)=\frac{2\Lambda(\varpi)^{-1}\mu(\varpi)^{-1}q^{s+1}}{q^{2}+1}.$
6.9.3. Local period for type VIb
Suppose that $\pi$ is of type VIb. We recall that $\dim(V_{\pi}^{K_{0}({\mathfrak{p}})})=\dim({\mathscr{B}}(T,\Lambda)^{K_{0}({\mathfrak{p}})})=1$; thus, by [33, Theorem 9.3], there exists a unique element $B^{{\rm VIb}}\in{\mathscr{B}}(T,\Lambda)^{K_{0}({\mathfrak{p}})}$ satisfying $B^{{\rm VIb}}(1_{4})=1$. We fix the datum $(\ell_{\pi},\xi^{\pi})$ by $\ell_{\pi}(\xi^{\pi})=B^{{\rm VIb}}$. The following result is immediate from Proposition 6.13.
Proposition 6.18.
Let $\pi$ be a smooth admissible irreducible unitary representation of type VIb. Suppose that $:F^{\times}\to{\mathbb{C}}^{1}$ is an unramified character. When $\phi=\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}$,
7. Computation of Gross-Prasad period for ramified characters
We shall add a new case to the local computations in [7, §2]. Let $F$, ${\mathfrak{o}}$, $T$, $L$ and $\Lambda$ be as in §6. Suppose that $T=\left[\begin{smallmatrix}{\mathsf{a}}&\mathsf{b}/2\\ \mathsf{b}/2&\mathsf{c}\end{smallmatrix}\right]\in{\rm Sym}_{2}(F)$, the character $\Lambda$ of $L^{\times}$ and an irreducible unitary representation $(\pi,V_{\pi})$ of $\mathsf{G}(F)/{\mathsf{Z}}(F)$ satisfy
(7.1) ${\mathsf{a}},\mathsf{c}\in{\mathfrak{o}}^{\times},\mathsf{b}\in{\mathfrak{o}},{\mathsf{d}}:=\mathsf{b}^{2}-4{\mathsf{a}}\mathsf{c}\in{\mathfrak{o}}^{\times}$, $L/F$ is an unramified field extension
(7.2) $\Lambda$ is of conductor $1+{\mathfrak{p}}_{L}$,
(7.3) $(\pi,V_{\pi})$ is $K$-spherical.
Let $(\cdot\mid\cdot)_{\pi}$ be the $\mathsf{G}(F)$-invariant hermitian inner product on $V_{\pi}$, and $\xi_{0}\in V_{\pi}$ be a $K$-fixed unit vector; define a ${\mathsf{Z}}(F)K$-bi-invariant function $\Phi_{\pi}$ on $\mathsf{G}(F)$ as
$\Phi_{\pi}(g):={(\pi(g)\xi_{0}\mid\xi_{0})_{\pi}},\quad g\in\mathsf{G}(F).$
Our main concern is the limit
$J_{0}^{T}(g):$ $=\lim_{k\to\infty}\int_{{\mathfrak{p}}^{-k}\mathsf{N}({\mathfrak{o}})}\Phi_{\pi}(b^{-1}ngb)\psi_{T}(n){\rm{d}}n$
(7.4) $=\lim_{k\to\infty}\int_{({\mathfrak{p}}^{-k})^{3}}\Phi_{\pi}\left(b^{-1}\left[\begin{smallmatrix}1&&x&y\\ &1&y&z\\ &&1&\\ &&&1\end{smallmatrix}\right]gb\right)\psi({\mathsf{a}}x+\mathsf{b}y+\mathsf{c}z){\rm{d}}x{\rm{d}}y{\rm{d}}z,\quad g\in\mathsf{G}(F),$
where $b:=h(0,1)=\left[\begin{smallmatrix}\varpi^{2}&&&\\ &\varpi&&\\ &&1&\\ &&&\varpi\end{smallmatrix}\right]\in\mathsf{G}(F)$. The limit (7.4) makes sense; Indeed, by [24], the values of integrals are independent of large enough $k$. The local period $J_{0}^{T,\Lambda}$ and its normalization $J^{T,\Lambda}$ are defined by
$J_{0}^{T,\Lambda}:=\int_{F^{\times}\backslash L^{\times}}J_{0}^{T}(\mathsf{m}(\tau,\det\tau))\Lambda(\tau){\rm{d}}\tau,\quad J^{T,\Lambda}:=\frac{L(1,\pi;{\rm Ad})L(1,\chi_{L/F})}{\zeta_{F}(2)\zeta_{F}(4)L(\frac{1}{2},\pi\times\mathcal{AI}(\Lambda^{-1}))}J_{0}^{T,\Lambda}.$
Note that, in our setting that $\pi$ is unramified and $\Lambda$ is ramified, the local $L$-factor $L(s,\pi\times\mathcal{AI}(\Lambda^{-1}))$ is identically $1$.
Theorem 7.1.
Under condition (7.1), we have $J^{T,\Lambda}=q^{-4}$.
For the rest of this section, we sketch a proof of this theorem, introducing key lemmas that are used in the proof. Let $v:F\to{\mathbb{Z}}\cup\{\infty\}$ be the normalized valuation defined by $|x|=q^{-v(x)}$. Recall the double coset decomposition $\mathsf{G}(F)=\sqcup_{\ell,m{\geqslant}0}{\mathsf{Z}}(F)K\cdot h(\ell,m)\cdot K$.
Lemma 7.2.
Let $x,y,z\in F$.
  • (i)
    If $\left[\begin{smallmatrix}1&&\varpi^{-2}x&\varpi^{-1}y\\ &1&\varpi^{-1}y&z\\ &&1&\\ &&&1\end{smallmatrix}\right]\in{\mathsf{Z}}(F)K\cdot h(l,m)\cdot K$, then
    $\ell$ $=2\{\min(0,-2+v(x),-1+v(y),v(z))$
    $\qquad-\min(0,-2+v(xz-y^{2}),-2+v(x),-1+v(y),v(z))\},$
    $m$ $=\min(0,-2+v(xz-y^{2}),-2+v(x),-1+v(y),v(z))$
    $\qquad-2\min(0,-2+v(x),-1+v(y),v(z)).$
  • (ii)
    If $\left[\begin{smallmatrix}\varpi&&\varpi^{-1}x&y\\ &\varpi^{-1}&\varpi^{-2}y&\varpi^{-1}z\\ &&\varpi^{-1}&\\ &&&\varpi\end{smallmatrix}\right]\in{\mathsf{Z}}(F)K\cdot h(l,m)\cdot K$, then
$\ell$ $=2\{\min(0,v(x),-1+v(y),v(z))$
$\qquad-\min(0,v(xz-y^{2}),v(x),1+v(y),v(z))\}+2,$
$m$ $=\min(0,v(xz-y^{2}),v(x),1+v(y),v(z))$
$\qquad-2\min(0,v(x),-1+v(y),v(z)).$